OpenStudy (anonymous):

divide 8 4/5 divided by 1 1/2

OpenStudy (anonymous):

@salemlover352

OpenStudy (trojanpoem):

\[a \frac{ b }{ c } = \frac{ ac + b }{ c }\] \[8\frac{ 4 }{ 5 } = \frac{ 44 }{ 5 }\] \[1\frac{ 1 }{2 } = \frac{ 3 }{ 2 }\] Division is equivalent to multiplying times the inverse: Inverse of a= 1/a 44/5 * 2/3 = 88/ 15

OpenStudy (anonymous):

so wats the answer